OpenStudy (anonymous):

solve by the substitution method 5x+7y=53 -5x+y=19 The answer I got is -4.25 Can anyone help with this

OpenStudy (anonymous):

Add the two equations together gets you 8y = 53 + 19 Then divide by 8 to isolate y by itself to get y =53+19/8 . you'll get 9 Then simply use one of the equations to solve for x

hero (hero):

Yeah, the answer is definitely not -4.25

hero (hero):

Do what outkast said

OpenStudy (anonymous):

When it asks for Substitution think of it as trying to cancel one of the variable (either x or y) and then solving for the leftover variable. Then take that number and plug it in to get the other variable. So, in your case, 5x + 7(9) = 53 = 5x +63 = 53 -63 -63 5x = -10 x = -2 i hope this helps

OpenStudy (anonymous):

would the ordered pair be (-2,9)?

OpenStudy (anonymous):

I believe so